OpenStudy (anonymous):

audible range of frequencies to which human ear responds varies between 20 to 20,000 Hz. express the range in terms of (1) wave lengths in air (2)time period and angular frequency. velocity of sound in air is 350 m/s.

OpenStudy (anonymous):

\[v=f \lambda \]and so \(\lambda =v/f \) time period is \[T =1/f \] Here, v=velocity of sound, lambda=wavelength of the sound waves f=frequency

OpenStudy (anonymous):

actually i had lil bit problem with the answer so i post this question but any thanks rajathsbhat

OpenStudy (anonymous):

cool. :)